### v3

# Kishu: Time-Traveling for Computational Notebooks

Abstract.
Computational notebooks (e.g., Jupyter, Google Colab) are widely used by data scientists. A key feature of notebooks is the interactive computing model of iteratively executing cells (i.e., a set of statements) and observing the result (e.g., model or plot). Unfortunately, existing notebook systems do not offer time-traveling to past states: when the user executes a cell, the notebook session state consisting of user-defined variables can be irreversibly modified—e.g., the user cannot ’un-drop’ a dataframe column. This is because, unlike DBMS, existing notebook systems do not keep track of the session state. Existing techniques for checkpointing and restoring session states, such as OS-level memory snapshot or application-level session dump, are insufficient: checkpointing can incur prohibitive storage costs and may fail, while restoration can only be inefficiently performed from scratch by fully loading checkpoint files.
In this paper, we introduce a new notebook system, Kishu, that offers time-traveling to and from arbitrary notebook states using an efficient and fault-tolerant incremental checkpoint and checkout mechanism. Kishu creates incremental checkpoints that are small and correctly preserve complex inter-variable dependencies at a novel Co-variable granularity. Then, to return to a previous state, Kishu accurately identifies the state difference between the current and target states to perform incremental checkout at sub-second latency with minimal data loading. Kishu is compatible with 146 object classes from popular data science libraries (e.g., Ray, Spark, PyTorch), and reduces checkpoint size and checkout time by up to 4.55$\times$ and 9.02$\times$, respectively, on a variety of notebooks.
1. Introduction
Computational notebooks (e.g., Jupyter (Jupyter, 2023; Team, 2023a), Rstudio (Posit Software, 2023)) are widely used by data scientists (Ormond, 2018; Perkel, 2018). A key feature of the notebook workflow is iterative code execution and result observation (Amershi et al., 2019; Chattopadhyay et al., 2020), which is highly compatible with the incremental nature of data science tasks, such as interactive tutorials (Johnson, 2020), data exploration (Crotty et al., 2015; Zgraggen et al., 2014; Dunne et al., 2012), visualization (Eichmann et al., 2020), and model tuning (Wagenmakers and Farrell, 2004; Bergstra and Bengio, 2012). This iterative workflow is enabled by notebooks systems being stateful—to do work, users would start a session, then as users execute code in the notebook system, the results are held in the session state as user-defined variables (e.g., loaded datasets, fitted models).
Limitation: no Time-Traveling for Notebooks
Oftentimes, during a workflow, users would like to revert changes made to the session state (i.e., time-travel), such as to undo a modification (e.g., restore a dropped column of a dataframe (StackOverflow, 2024)), restoring an overwritten variable (Groups, 2024), or perform reverse debugging (Brachmann and Spoth, 2020). Unfortunately, unlike program debuggers (e.g., gdb) (Phang et al., 2013; Barr and Marron, 2014; GDB, 2024), relational databases (e.g., PITR in PostgreSQL and MySQL (Group, 1996a; Oracle, 2024b)) or interactive data systems (Dunne et al., 2012; Crotty et al., 2015; Kraska, 2021) which support time-traveling to past program states, existing notebook systems do not natively keep track of past session states: cell executions cannot be undone, e.g., the user cannot ’un-drop’ a dataframe column. If the user executes a cell that alters the session state, a common approach to restore the previous state would be to restart the kernel and then (painstakingly) re-run past cells in the correct order. While code versions can be saved using tools such as Git (Git, 2024b) or native commands (e.g., Jupyter’s %checkpoint (Team, 2023b) [^1] Despite its name, %checkpoint only stores cell code and not objects in the state.) to simplify identifying cells to rerun for restoration, cell reruns can still be time-consuming (e.g., re-training an ML model) and/or result in incorrect restoration (e.g., random train-test splits). Another approach is for the user to periodically checkpoint the session state (e.g., memory dump (CRIU, 2023; Ansel et al., 2009) or session state serialization (Foundation, 2023h; Weinman et al., 2021)) to storage or a managed database (e.g., KV-store (Team, 2023c)). Then, users can load an appropriate checkpoint file to restore the session state. However, performing session checkpointing and restoration using these tools is limiting: checkpointing can incur prohibitive costs (§ 7.3, § 7.4) and may fail on certain workloads (e.g., GPU (cri, [n.d.])), and restoration can either (1) only be (inefficiently) performed from scratch, requiring completely loading a checkpoint file (Foundation, 2023h) and/or killing the current kernel (CRIU, 2023), or (2) may be incorrect, breaking inter-variable relations (Team, 2023c).
Our Goal: Generalizable, Correct, and Efficient Time-Traveling
We propose Kishu, a notebook system that enables time-traveling between session states: as the user executes cells, Kishu tracks the session state evolution while writing per-cell incremental checkpoints containing differing data between successive states (i.e., the state delta) for returning to any past state via an incremental checkout later. Kishu pursues three challenging goals—Delta-Efficient Checkpoint: Kishu aims to minimize incremental checkpointing overhead by exploiting the small per-cell deltas typical of data science workflows (§ 2.2), but also avoid high detection overhead in the face of complex access patterns and inter-variable dependencies. Correct & Non-intrusive Checkout: Kishu aims to restore past states in the same session non-intrusively by leveraging existing objects in the kernel (that don’t need updating) to minimize data loading costs, while still guaranteeing checkout accuracy as if it completely loaded a checkpoint file. Generalizability: Kishu aims to support checkpointing/checkout for almost all notebook libraries (and/or use cases), of which there is a large variety, e.g., notebooks can perform distributed computing (e.g., Spark (Zaharia et al., 2010)) or move data off-CPU (e.g., GPUs (Foundation, 2024a)). If Kishu can achieve these goals, Kishu will allow users to undo almost any executed cell that undesirably modifies the state as if it never occurred by quickly checking out to the pre-execution state at the cost of minimal workflow overhead.
Our Approach
Our core idea for achieving our goals is to capture the state delta with low overhead, but at sufficiently high granularity using information exclusive to the application level, as follows:
First, for delta-efficient incremental checkpointing, Kishu utilizes low-overhead live analysis (e.g., namespace patching) to track session state evolution at a novel Co-variable granularity (i.e., connected components of objects). Then, Kishu writes and versions Co-variables with the checkpoint graph representing the user workflow in terms of cell executions to minimize delta storage overhead.
Second, for correct incremental checkout, Kishu identifies the difference between the current and target states at the aforementioned Co-variable granularity according to the checkpoint graph. Then, it replaces (only) Co-variables that need updating in the state by loading data from appropriate incremental checkpoints, minimizing data load time for checkout and transparently restoring the state in the same kernel process without interruption, at sub-second latency.
Third, Kishu achieves generalizability and fault-tolerance through fallback recomputation. If a Co-variable cannot be stored in a checkpoint (e.g., it contains an unserializable object such as a hash (Foundation, 2023c)) or fails to load upon checkout, Kishu can efficiently reconstruct it upon checkout via finding the shortest path combining intermediate data loading and cell re-running according to the checkpoint graph.
Difference from Existing Work (LABEL:tbl:existing_work)
Our work enables efficient time-traveling for computational notebooks through significantly different techniques vs. existing work. OS-level tools (CRIU, 2023; Ansel et al., 2009) can incrementally checkpoint notebook states, but fail to exploit the fine-grained deltas of data science workflows (§ 2.2), cannot incrementally restore, and fail on remote objects (e.g. Ray (Moritz et al., 2018), on-device data (PyTorch, 2024)). Existing application-level tools for saving state (Foundation, 2023g, f; Li et al., 2023a; Weinman et al., 2021) lack both the incremental checkpointing and restoration capabilities of our work. Works for versioning notebook cell code (Brachmann et al., 2019; Guo and Seltzer, 2012; Team, 2023b) help identify cells to rerun for restoration but do not directly enable it. Incremental checkpointing/PITR in DBMS (Antonopoulos et al., 2019; Oracle, 2024a)/HPC (Lin et al., 2020; Keller and Bautista-Gomez, 2019) and time-traveling DBMS (Soroush and Balazinska, 2013; Schule et al., 2019) focus on robust and fast logging of table/KV-store updates, while our work focuses on delta computation for complex, interdependent objects unique to notebooks, i.e., what to log (§ 2.2). Orthogonal works include notebook systems for speeding up data exploration using non-time-travel methods (e.g., code recommendation (Lee et al., 2021; Li et al., 2023b)).
Contributions
According to our motivations in § 2, we implement Kishu (§ 3), a notebook system with the following contributions:
  • State Delta Detection. We introduce our modeling of session state evolution at a novel Co-variable granularity, and our correct and efficient delta detection at this granularity. (§ 4)
  • State Versioning. We introduce our delta-based session state versioning with the Checkpoint Graph, which enables efficient and fault-tolerant incremental checkpointing and checkout. (§ 5)
  • Time-traveling. We show via experimental evaluation that Kishu’s time-traveling is compatible with 146 classes from popular data science libraries and reduces checkpoint size and checkout time by up to 4.55$\times$ and 9.02$\times$, respectively. (§ 7)
2. Motivation
This section describes use cases for time-traveling notebooks (§ 2.1), characteristics of notebook workloads (§ 2.2), and, accordingly, our intuition for time-traveling (§ 2.3) and capturing state delta (§ 2.4).
2.1. Why is Time Traveling Useful?
Time-traveling computational notebooks can enable users to efficiently undo cell executions and perform path-based exploration.
Undoing Cell Executions
Data operations can be irreversible (e.g., df=df.drop_col(’a’)), and users may want to return to the previous state if the operation outcome is undesirable (Groups, 2024; Brachmann and Spoth, 2020). To enable interruption-free time-traveling, we can checkpoint the state delta to storage after each operation such that the session state prior to performing the operation can be returned to via loading the appropriate deltas. We empirically study this use case in § 7.5.1.
Path-based exploration
Data scientists often manually create branching, out-of-order cells, where each intended execution path consists of only a subset of notebook cells (e.g., cleaning steps for different models) (Rule et al., 2018; Kery et al., 2018). If we can efficiently persist all variations of objects in different paths as incremental deltas w.r.t. a shared state, users can efficiently switch paths for comparisons: only the (small amount of) data differing between paths need to be updated via loading deltas. We empirically study this use case in § 7.5.2.
2.2. Characteristics of Notebook Workloads
Data scientists commonly use computational notebooks for exploratory work (Kery et al., 2018), which exhibit these key characteristics:
Think Time
Notebook workflows often follow a loop of users sequentially [^2] Existing notebook systems, e.g., Jupyter, do not support concurrent cell executions. writing and running cell code, then observing output (Rule et al., 2018). The notebook kernel can be inactive mid-loop—$\sim$10 seconds of think time (Eichmann et al., 2020) when users decide what cells to write and run next. Notebook systems use think time for computations (e.g., speculative materialization (Xin et al., 2021)). Kishu can also leverage think time for incremental checkpointing.
Complex Access Patterns
Python functions in notebooks can exhibit complex access patterns (e.g., accessing non-parameter variables), making it difficult and/or costly to analyze changes. Hence, some works use a rule-based approach to hard-code effects of common functions (e.g., print not modifying the namespace) (Macke et al., 2020); however, as data scientists also commonly implement and import custom functions into notebooks (Rule et al., 2018), an efficient approach that handles arbitrary functions is highly desired.
Incremental Operations
Users often run cells incrementally, each consisting of only a few lines of code accessing a small number of variables, for fast iteration (Kery et al., 2018). Accordingly, we observe in on of our test notebooks, Sklearn (LABEL:fig:background_workload_characteristics, top), that 40/44 cells access ¡10% of the state data. For these workloads, Kishu’s incremental checkpointing can offer larger benefits.
Data Modifications
Notebook providers often limit session memory consumption (Google, 2024; Kaggle, 2024). Hence, in addition to creating new data, users also modify/delete existing data to conserve memory (e.g., after they have been used to generate relevant cell outputs). This is seen in Sklearn (LABEL:fig:background_workload_characteristics, bottom), where updated data is, on average, split 45%:55% between creations and modifications. Kishu must efficiently undo these in-place updates during checkout.
2.3. Enabling Time Traveling
We discuss the pros and cons of incremental checkpointing and checkout approaches for enabling time-traveling to a previous state.
OS-level Memory Snapshots
Tools such as CRIU (CRIU, 2023) can snapshot notebook process memory to save session state data. (Incremental Checkpointing) Subsequent snapshots can be made incrementally w.r.t. previous snapshots storing only dirty memory pages. However, memory-page granularity is too coarse for incrementally checkpointing notebook workloads as Python data structures (e.g., lists) can be fragmented, leading to simple operations (e.g., in-place sorting a list) causing multiple dirty pages and high checkpoint costs (§ 7.3). (Complete Checkout) Memory snapshots must be entirely loaded to restore a state and additionally require killing the existing notebook process (to avoid PID conflicts) before restoration, which is not seamless and incurs high data loading costs (§ 7.5). Furthermore, OS-level snapshotting is limited to single processes, hence fail on notebooks utilizing multiple/remote processors (§ 7.3).
Application-level Incremental Checkpoint and Checkout (Ours)
We can utilize (1) application-level information to detect state deltas at a finer granularity compared to memory snapshots and (2) existing data within the kernel for incremental checkout to address the drawbacks of OS-level snapshotting: (Incremental and Generalized Checkpointing) We track state deltas at a more logical Co-variable granularity (described and motivated in the next session) by tracking references in the notebook for incremental checkpointing. For generalizability, we utilize an object’s reduction (i.e., __reduce__ (Foundation, 2023f)) as storage instructions to checkpoint multiprocessing and off-CPU workloads (§ 6.1). (Incremental Checkout) As we have direct access to the notebook kernel at the application level, we can incrementally checkout by only updating data in the kernel that differ between the current and target states after computing said difference (e.g.., via state snapshots (Bernstein and Goodman, 1983)) (§ 5.2).
2.4. Tracking State Delta for Time Traveling
We discuss pros and cons of methods of tracking the state delta at the application level for incremental checkpointing and checkout.
Provenance-based Tracking
Some existing notebook systems (Macke et al., 2020; Koop and Patel, 2017; Li et al., 2023a) track session state evolution via provenance-based code analysis (e.g., via ASTs (Foundation, 2023a)) at a variable-level granularity. Pure static code analysis requirew conservativeness on identifying changed data w.r.t. control flows (e.g., if (x<1):) and external function calls (§ 2.2) causing false positives (e.g., by assuming an untaken branch as taken) and large deltas; hence, these systems augment code analysis with varying degrees of live instrumentation at cell runtime (e.g., resolving the value of x when evaluating the aforementioned if statement) (Macke et al., 2020), which can potentially result in high overhead (e.g., repeated resolutions in looping control flows, § 7.6).
Co-variable Granularity Live Tracking (Ours)
In order to avoid potential efficiency issues caused by the runtime resolution of provenance tracking, we propose to perform live object comparison (i.e., comparing data before and after execution) only after each cell execution to track per-cell updates. Our intuition is that while doing so at variable-level granularity (i.e., like existing provenance trackers) by comparing all objects in the state can be expensive (§ 7.6), it is also unnecessary as storing/loading individual variables (e.g., with variable-level KV-stores (shove, 2024; 0xnurl, 2024)) during checkpointing/checkout risks breaking shared references (Li et al., 2023a). We hence track updates at a coarser Co-variable level—connected components of objects (w.r.t. pointer references): we can reason based on access patterns which Co-variables were possibly/definitely not updated by each cell to limit object comparisons, and correctly store/load Co-variables as if they are independent data tables. LABEL:fig:background_variable_blob depicts our idea: {ser,obj} is a Co-variable (red) as the reachable objects of the variables overlap (&ser[1] = &obj.foo). {df} is another Co-variable (blue), and users cannot reach objects under df from objects under ser or obj via references. Notably, Co-variables are the minimum granularity which data in the state can be stored/loaded without risking breaking shared references. § 4 formally describes Co-variables and how we correctly and efficiently capture state delta at this granularity.
Motivating Example (LABEL:fig:background_example)
Suppose a data analyst is performing text mining. They load the corpus (Cell 1), define category lists (Cell 2), and sort texts by sentiment into the lists (Cell 3). The analyst incrementally checkpoints the state after each cell execution. (1) Incremental Checkpointing: The analyst tests a mapping to clean the contained text in the lists on sad_ls (Cell 4, blue). Due to its interleaved construction (with other lists), the list sad_ls is fragmented; incrementally checkpointing the state at memory page granularity for Cell 4 (w.r.t. Cell 3) results in all pages overlapping with sad_ls being copied. However, a Co-variable granularity incremental checkpoint stores only (the bytestring of) sad_ls.
(2) Incremental Checkout: The analyst undoes the mapping in Cell 4 due to poor results. Returning to Cell 3’s state by (completely) loading a memory snapshot is slow as it requires reloading the corpus. However, by identifying that the states of Cells 3 and 4 differ only by sad_ls, we can only load the value of sad_ls from Cell 3 (red) to replace value of sad_ls from Cell 4 (blue) to perform incremental checkout without touching the rest of the session state.
3. System Overview
This section presents Kishu components (§ 3.1) and workflow (§ 3.2).
3.1. Kishu Components
Kishu (see LABEL:fig:system_overview) interacts with notebook sessions through inserting non-intrusive hooks. These hooks allow Kishu to transparently (1) monitor the kernel namespace to track session state evolution, (2) write data in the session state to storage for incremental checkpointing, and (3) alter the session state when checkout is requested.
Patched Namespace
On session start, Kishu patches the namespace to monitor accesses to its contents during cell executions (§ 4.3). It tracks user-referenced variable names to identify candidate Co-variables to check for updates in, which are passed to the Delta Detector to compute the Co-variable granularity state delta.
Delta Detector
The Delta Detector computes the state delta based on the candidates identified from the Patched Namespace (i.e., which of the candidate Co-variables were actually updated by the cell execution). We discuss the Kishu’s delta detection in (§ 4).
Checkpoint Graph
The Checkpoint Graph is a tree-like structure analogous to Git’s commit graph (Git, 2024a), in which Kishu writes, stores, and versions incremental checkpoints consisting of the updated Co-variables (i.e., the state delta) of each cell execution (§ 5.1). The incremental checkpoints stored in the Checkpoint Graph are used by the State Loader to perform incremental checkout.
State Loader
The State Loader restores to a session state upon requested checkout. It first identifies the difference between the current session state (i.e., existing items in the namespace) and the target state according to the Checkpoint Graph, then loads only the necessary data from the Checkpoint Graph for replacing the Co-variables that need updating (§ 5.2).
Data Restorer
The Data Restorer is a mechanism that utilizes fallback recomputation to restore missing data for checkout (e.g., Kishu failed to serialize the data during prior checkpointing). It reconstructs missing data by combining loading dependent data and cell re-runs according to the Checkpoint Graph. (§ 5.3)
3.2. Kishu Workflow
This section describes Kishu’s operations during a notebook workflow. Users will interact with Kishu using the in-Jupyter Command Palette (LABEL:fig:system_overview), such as to attach it to a started notebook session; [^3] Kishu’s interface is similar to that of ElasticNotebook; a depiction can be found in its demo paper (Li et al., 2024). However, the core functionality and techniques used are different. Kishu then monitors the namespace for state deltas to incrementally checkpoint after each cell execution (if automatic checkpointing is enabled), and perform checkout to a previous state when requested.
Attaching Kishu to a Notebook Session
When initializing a notebook session, init attaches Kishu to the kernel, which will patch the namespace and initialize the Checkpoint Graph on storage.
Incremental Checkpointing
After each cell execution, the Delta Detector uses the Patched Namespace to identify updated Co-variables and stores them in a new incremental checkpoint/node with a unique checkpoint_id on the Checkpoint Graph. The user may view the stored graph, checkpoints, and their IDs with log.
Incrementally Restoring a State
Kishu will restore a previous session state with checkout checkpoint_id. The State Restorer identifies differing Co-variables (i.e., to restore) between the current and target states according to the Checkpoint Graph, then accordingly loads only the necessary data for restoration. If necessary, the Data Restorer reconstructs data that is missing (i.e., failed to serialize) or failed to load via fallback recomputation.
4. Accurate and Fast Delta Detection
This section describes how Kishu correctly detects Co-variable granularity state deltas necessary for incremental checkpointing and checkout (§ 2.4). We formally describe Co-variables in § 4.1 and how we detect Co-variable updates correctly (§ 4.2) efficiently (§ 4.3).
4.1. Co-variables
In this section, we introduce the Co-variable—Kishu’s granularity for efficient incremental checkpointing and checkout.
Preliminary: Variables, Objects and Reachability
In Python and the Jupyter Notebook ecosystem, variables and objects are 2 distinct concepts: A variable is a named entity from which various objects are reachable—for an example list ls=[1,2,3], the list name (ls) is a variable and each list element (1, 2, 3) is an object. We define reachability reference-wise, i.e., object y is reachable from variable x if y can be accessed from x through a chain of references. Some common reachability patterns include subscripting (e.g., y = x[0]), class member (e.g., y = x.attr), and attribution (e.g., y = x.__dict__). Given our distinction between variables and objects and reachability definition, we now define Co-variables as follows:
Definition 0.
A Co-variable is a set of variable names $\mathcal{X}=\{x_{1},...,x_{i}\}$ from which the reachable objects form a maximally connected component. That is, for any variable $y$ not in the set, the objects reachable from $x_{1},...,x_{i}$ are not reachable from $y$.
A Co-variable can consist of one name (e.g., a primitive, x=1) or multiple names from which same objects can be reached (i.e., shared references). For example, in LABEL:fig:checkpoint_variable_update, the string object ’b’ is reachable from both Pandas Series ser and object obj via subscript and class member respectively, hence {ser,obj} is a Co-variable. Co-variables are self-contained by definition, i.e., there are no inter-Co-variable references. They can be modified by cell executions:
Definition 0.
A Co-variable $\mathcal{X}=\{x_{1},...,x_{i}\}$ is modified by a cell execution if the graph structure of the connected component of objects reachable from $x_{1},...,x_{i}$ is modified, counting both node (i.e., object) and edge (i.e., reference) additions and deletions.
For example, the Co-variable {ser,obj} in LABEL:fig:checkpoint_variable_update is modified node-wise with an in-place update “ser.replace’’ (bottom), and edge-wise with a re-assignment “obj.foo = ser[2](bottom-right). Co-variables can also be created and deleted via split and merge (right): the Co-variable {obj,ser} is deleted via a split as obj and ser longer share references, and the Co-variable {obj, st} is created through a merge. We collectively refer to Co-variable modifications, creations, and deletions as updates; the set of Co-variables updated by a cell execution is the execution’s state delta.
4.2. Accurate State Delta Detection
This section describes how Kishu accurately detects Co-variable membership (i.e., which variables form a Co-variable) and updates.
VarGraphs
Kishu uses VarGraphs to detect Co-variable membership and updates. The VarGraph is a graph structure constructed from each variable in the namespace that captures its reachable objects. An example is shown in LABEL:fig:checkpoint_id_graph: each node in a variable’s VarGraph corresponds to a reachable object, containing the (1) object type, (2) memory address, and one of (3) pointers to other reachable objects (i.e., children) for non-primitive objects, or (4) value for primitive objects. For example, the node for the list reachable from ser contains 3 child pointers to the 3 nodes for strings ’a’, ’b’, and ’c’, and the node for string ’b’ holds its value ’b’.  [^4] The VarGraph is inspired by ElasticNotebook’s ID graph (Li et al., 2023a) which captures reachable objects’ memory addresses; VarGraphs uniquely contain datatypes and primitive values for additional robustness (e.g., detecting a different primitive in the same address).
Detecting Co-variable membership
Co-variable membership is determined by intersecting VarGraphs. For example, in LABEL:fig:checkpoint_id_graph, ser and obj form a Co-variable as the node ’b’ is in both graphs (red).
Detecting Co-variable updates
Co-variable updates is determined by comparing VarGraphs before and after cell executions. A graph structure modification and/or a node attribute change (e.g., object memory address or type) indicates an update to the Co-variable.
Accuracy Guarantee
As Kishu constructs VarGraphs following object rechability, it detects Co-variable updates with no false negatives (verified empirically in § 7.2.1). However, Kishu’s update detection is conservative: there may be false positives if objects are dynamically generated (e.g., datatype objects) with a different memory address each time during VarGraph construction/object traversal, or cannot be traversed into (i.e., lacking referencing instructions, e.g., generators (Foundation, 2023b), which Kishu assumes to be updated on access).
4.3. Efficient State Delta Detection
This section describes how Kishu speeds up Co-variable update detection. Identifying updates across the entire global namespace via VarGraphs can be expensive (due to object traversals); Kishu needs to reduce the number of Co-variables (i.e., portion of namespace) it checks after cell executions without reducing detection accuracy.
Identifying Possibly Updated Co-variables
Cell executions in Jupyter Notebook interact with the global namespace (i.e., globals()). Therefore, if Kishu can capture variable references in the cell execution, it can reason about which Co-variables were possibly updated (and which ones were definitely not), as follows:
Definition 0.
A Co-variable $\mathcal{X}=\{x_{1},...,x_{i}\}$ is accessed by a cell execution if any variable $x_{1},...,x_{i}$ is accessed (via getting, setting, or deletion) during the cell execution.
Co-variable accesses indicates possible updates (e.g., via a member function call ser.replace). Kishu patches the accessor, setter, and deletion methods of the global namespace (LABEL:fig:checkpoint_detect_fast) to capture variable (hence Co-variable) accesses, which helps identify the possibly updated Co-variables: if the members of a Co-variable $(x_{1},...,x_{i})$ overlaps with the cell execution’s accessed variables, then it may have been updated (e.g., {ser,obj} in LABEL:fig:checkpoint_detect_fast): Kishu will verify the update by (1) re-generating VarGraphs for its member variables, (2) comparing the VarGraphs with those before the cell execution to identify modifications, and (3) intersecting the VarGraphs amongst variables of accessed Co-variables to identify merges and splits. Otherwise, the Co-variable surely wasn’t updated and Kishu skips its check for this cell execution (e.g., {df} in LABEL:fig:checkpoint_detect_fast, greyed out).
Lemma 0.
A Co-variable $\mathcal{X}=\{x_{1},...,x_{i}\}$ can be updated by a cell execution only if at least one of $x_{1},...,x_{i}$ was accessed in the code.
Proof.
Suppose not, i.e., Co-variable $\mathcal{X}$ has empty intersection with the accessed variables and was updated. Then, the Co-variable must have updated through another variable $y$ that was not part of the Co-variable $x_{1},...,x_{i}$ before the start of the cell execution. Due to Co-variables’ self-containment (§ 4.1), the user cannot possibly access objects reachable from $x_{1},...,x_{i}$ via $y$ during the cell execution without creating a reference by using one of $x_{1},...,x_{i}$ first (e.g., $y.foo$ = $x\_i$), but doing so violates our assumption. ∎
As only a small portion of variables are accessed per cell in a typical data science notebook, Kishu significantly reduces delta detection overhead with this approach (empirically verified in § 7.6).
Remark
As Kishu patches the global namespace of the notebook session, it is impossible for the user to use variables from within the notebook (i.e., to modify objects) undetected. Hence, Kishu will not misidentify Co-variables possibly updated via references.
The user may still use non-referencing methods to update data such as with direct C-pointer-based modifications, but these cases are rare in notebooks (found in 0 out of 60 surveyed notebooks (Li et al., 2023a)). Some libraries, such as NumPy (Team, 2024a), do perform memory-based updates (e.g., via slicing). However, the objects are supported by Kishu as their memory-based updates are still invoked via referencing (e.g., arr[0,1] += 1, empirically verified in § 7.2.1).
5. Inc. Checkpoint & Checkout
This section describes Kishu’s efficient time-traveling with the Co-variable granularity state deltas. We describe Kishu’ incremental checkpointing in § 5.1, Kishu’s incremental checkout in § 5.2, and how Kishu time-travels to and from notebook states with problematic (e.g., unserializable) data in a fault-tolerant manner in § 5.3.
5.1. Incremental Checkpointing
This section describes how Kishu performs incremental checkpointing by writing and managing per-cell-execution checkpoints containing the updated Co-variables with the Checkpoint Graph.
Checkpoint Graph
The Checkpoint Graph is a directed tree of (incremental) checkpoints representing the branch-based state evolution. Nodes are added for each checkpoint Kishu performs, and timestamped according to the completion time $t$ of the corresponding cell execution (we refer to the timestamped node and cell execution as node $t$ and CE $t$, respectively, for simplicity). The Checkpoint Graph maintains a head node, which tracks the user’s current state. Each node $t$ contains the state delta consisting of (only) Co-variables updated by CE $t$. Co-variables stored in the Checkpoint Graph are versioned according to their corresponding CE $t$:
Definition 0.
A Versioned Co-variable is a Co-variable-timestamp pair $(\mathcal{X},t)$ representing the Co-variable $\mathcal{X}$ updated by CE $t$.
Versioned Co-variables are analogous to versioned datasets: the same Co-variable (w.r.t. variable membership, i.e., $\mathcal{X}=\{x_{1},...,x_{i}\}$) can take on multiple values throughout a notebook session being updated by different cell executions. LABEL:fig:checkpoint_graph show an example: CE $t_{3}$ creates the Co-variable {plot}, hence it is stored in node $t_{3}$ (red) as the Versioned Co-variable $((plot),t_{3})$.
Writing into the Checkpoint Graph
Kishu creates a new node $t$ in the Checkpoint Graph after each CE $t$ containing (1) the versioned Co-variables in the state delta of CE $t$, (2) the code of CE $t$, and (3) the versioned Co-variables (possibly stored in previous checkpoints) accessed by CE $t$ (§ 4.3). For example, node $t_{3}$ in LABEL:fig:checkpoint_graph (blue) contains the code of CE $t_{3}$ (“plot=gmm.result()“) and its dependency on the versioned Co-variable from node $t_{2}$ $((gmm),t_{2})$ (dashed line). Notably, the state delta, code, and variable accesses are respectively analogous to the update, operation, and dependencies in database logging and versioning. The new node $t$ is written under the head node $s$, and and a parent-child relationship is added from the head node $s$ to the new node $t$ (now the new head node).
Handling Unserializable Data
If Kishu cannot write an updated Co-variable in the state delta into the Checkpoint Graph (e.g., it contains an unserializable object such as a generator (Foundation, 2023b) or hash (Foundation, 2023c)), Kishu simply skips its storage. Instead, upon checkout, the missing (unserializable) Co-variable will be restored through fallback recomputation enabled by the cell code and dependencies stored in the Checkpoint Graph node, which we discuss in § 5.3.
5.2. Efficient State Restoration
When (incremental) checkout is requested, Kishu aims to accurately restore the current state to the target state in the fastest manner possible. To do so, it must identify the contents of the target state according to its timestamp, analogous to how timestamped state snapshots are defined in MVCC (Bernstein and Goodman, 1983); instead of versioned tables, we identify Versioned Co-variables in the (timestamped) target state:
Definition 0.
The Session State [^5] Kishu stores snapshots of Session State metadata (i.e., references to contained Co-variables) in Checkpoint Graph nodes. at timestamp $t$ is a set of $n$ Versioned Co-variables $\{(\mathcal{X}_{1},t_{1}),...,(\mathcal{X}_{n},t_{n})\}$ such that for each $(\mathcal{X}_{j},t_{j})$, $1\leq i\leq n$:
  1. 1.
    $t_{j}$ is an ancestor of $t$ on the Checkpoint Graph.
  2. 2.
    There must not exist another versioned Co-variable $(\mathcal{Y}_{k},t_{k})$ such that $\mathcal{X}_{j}\cup\mathcal{Y}_{k}\neq\emptyset$ and $t_{k}$ is a child of $t_{j}$ and ancestor of $t$.
The session state at timestamp $t$ (state $t$ for brevity) is the set of all Versioned Co-variables that are (1) available in the namespace after CE $t$ and (2) have not been overwritten by a newer Versioned Co-variable prior to CE $t$. For example, in LABEL:fig:restore_diff, state $t_{3}$ (top-left) consists of Versioned Co-variables ($(\{plot\},t_{3})$, ($(\{gmm\},t_{2})$, and ($(\{df\},t_{1})$. It does not contain ($(\{gmm\},t_{1})$ as it was overwritten by CE $t_{2}$ (gmm.fit(k=3)) which writes ($(\{gmm\},t_{2})$. Each state $t$ dictates which Versioned Co-variables should be loaded from various nodes on the Checkpoint Graph for checkouts; to efficiently accomplish (incremental) checkout, Kishu identifies the difference between the current and target states in terms of the (versioned) Co-variables that need to be updated: some Co-variables do not need to be updated when converting the current state to the target state. They can be identified via the Checkpoint Graph:
Definition 0.
A Co-variable $\mathcal{X}$ is identical between the current state $t_{a}$ and target state $t_{b}$ if a Versioned Co-variable $(\mathcal{X},t_{c})$ exists in the session states of $t_{a}$, $t_{b}$, and $t_{c}$, where $t_{c}$ is the lowest common ancestor of node $t_{a}$ and node $t_{b}$. Otherwise, if no such $(\mathcal{X},t_{c})$ exists, then the Co-variable $\mathcal{X}$ has diverged between $t_{a}$ and $t_{b}$.
A Co-variable $\mathcal{X}$ is identical between current and target states $t_{a}$ and $t_{b}$ if its versioned counterpart is consistent across $t_{a}$, $t_{b}$, and $t_{c}$, i.e., no cell execution between (1) node $t_{a}$ and node $t_{c}$ and (2) node $t_{b}$ and node $t_{c}$ updated $\mathcal{X}$, hence does not need updating when checking out from $t_{a}$ to $t_{b}$. For example, in LABEL:fig:restore_diff, if checking out from current state $t_{5}$ to target state $t_{3}$, the Co-variable {df} (blue) is identical between the states as no cell execution between (1) node $t_{1}$ and node $t_{3}$ and (2) node $t_{1}$ and node $t_{5}$ updated it.
Otherwise, if the Co-variable $\mathcal{X}$ has diverged between the current and target session states, it will need to be updated (by either loading the appropriate Versioned Co-variable or deleting it) to complete the checkout to the target state. For example, the Co-variable {gmm} (red) has diverged between nodes $t_{5}$ and $t_{3}$ as their parents ($t_{4}$ and $t_{2}$) both updated gmm with their cell execution (fitting with k=3 and k=10), hence, gmm (and plot) needs to be updated via loading ({gmm},$t_{2}$) if checking out from state $t_{5}$ to state $t_{3}$.
Performing State Checkout
When checking out to the state at node $t$, The State Restorer (§ 3.1) performs the following steps:
  1. 1.
    Load the appropriate Versioned Co-variables from nodes (i.e., node $t$ and ancestors of node $t$) to update diverged Co-variables between the state of the current head node $s$ and node $t$.
  2. 2.
    Update/re-generate VarGraphs (§ 4.2) for updated Co-variables.
  3. 3.
    Move the head from node $s$ to the checked out node $t$.
Notably, the next cell execution will create a node in a new branch rooted at $t$ in the Checkpoint Graph, e.g., the graph in LABEL:fig:restore_diff is generated through the sequence $t_{1}\rightarrow t_{2}\rightarrow t_{3}\rightarrow(\text{checkout to }\,t_{1})\rightarrow t_{4}\rightarrow t_{5}$. If during checkout, a required Versioned Co-variable is missing (i.e., due to serialization failure, § 5.1) or fails to load (i.e., deserialization failure), Kishu restores it via fallback recomputation.
5.3. Robust Restoration
In this section, we describe how Kishu restores problematic data to achieve generalizable and fault-tolerant incremental checkout.
Fallback Recomputation
As each Checkpoint Graph node $t$ contains the code of CE $t$ and (2) which Versioned Co-variables $(\mathcal{X}_{j},t_{j})$ (stored in previous deltas, $t_{j}<t$) CE $t$ accessed (§ 5.1), any Versioned Co-variable $(\mathcal{X},t)$ in the state delta of node $t$ can be recomputed by (1) loading the accessed Versioned Co-variables from previous checkpoints into a temporary namespace, then re-running CE $t$. For example, in LABEL:fig:restore_fallback, suppose Versioned Co-variable ({plot}, $t_{3}$) (green) fails to load when checking out to session state $t_{3}$. The versioned Co-variable (({gmm}, $t_{2}$) is required to rerun CE $t_{3}$ (red); therefore, ({gmm}, $t_{2}$) is loaded from the parent node $t_{2}$, and after rerunning CE $t_{3}$ on the input ({gmm}, $t_{2}$), ({plot}, $t_{3}$) is restored.
Dynamic and Recursive Fallbacks
Kishu’s fallback recomputation is dynamic and recursive—if another Co-variable is missing or fails to load when retrieving the inputs (i.e., accessed Co-variables) for performing fallback recomputation, fallback recomputation can be recursively performed for that Co-variable. For example, if ({gmm}, $t_{2}$) from node $t_{2}$ fails to load (as part of fallback recomputation for ({plot}, $t_{3}$) from node $t_{3}$), it itself can be recomputed by loading ({gmm}, $t_{1}$) from node $t_{1}$ and rerunning CE $t_{2}$ (blue).
Remark
Kishu guarantees exact restoration for all serializable Co-variables during checkout, i.e., they will have the same bytestring representation before and after checkout if there are no hidden serialization errors (§ 6.2). While Kishu is capable of restoring problematic Co-variables via fallback recomputation, it currently does not support exactly restoring Co-variables that both (1) fail to store or load and (2) are created through non-deterministic means (e.g., random generators). This is a limitation similarly present in Spark (Zaharia et al., 2010) and Ray’s (Moritz et al., 2018) lineage-based fault tolerance; however, in our case, unserializable objects are sufficiently rare in data science libraries (§ 7.2), hence we consider this limitation to be acceptable.
6. Implementation and Discussion
This section describes Kishu’s implementation details (§ 6.1) and design considerations (§ 6.2).
6.1. Implementation
Integrating with Jupyter
Kishu is implemented as a separate application from the notebook process, usable without altering the base Jupyter application. Upon session initialization, Kishu places hooks into the kernel (pre_run_cell and post_run_cell (IPython, 2024b)) and patches the namespace (user_ns (IPython, 2024a)) (§ 3.1), allowing the standalone Kishu process to detect state deltas, write data to storage, and overwrite data in the namespace upon checkout transparently.
Serialization Protocol
The Pickle protocol (i.e., __reduce__ (Foundation, 2023f)) is used for (1) object serialization and (2) constructing VarGraphs for identifying Co-variables, i.e., object y is reachable from another object x if pickle(x) includes y. As Pickle is the de-facto standard (in Python) observed by almost all libraries (e.g., NumPy, PyTorch (et al., 2018)), Kishu can be used in almost all cases. Kishu’s per-Co-variable storage also enables mixing and matching serialization libraries for coverage: Currently, Kishu will try CloudPickle (cloudpipe, 2024) first, then use Dill (Foundation, 2023g) as a fallback for Co-variables that CloudPickle fails on.
Storing Checkpoints
Currently, Kishu uses SQLite (SQLite, 2024) to store Versioned Co-variables in the Checkpoint Graph. However, any storage mechanism can be used in its place—even in-memory ones if the user wants to maximize checkpointing/checkout efficiency.
6.2. Design Considerations
Silent Serialization Errors
Certain object classes may contain incorrect serialization instructions, which, despite being able to be stored/loaded to/from storage, result in silent errors. Kishu currently assumes that the serialization instructions are correctly implemented for all objects in terms of equality before and after pickling (i.e., x == unpickle(pickle(x)) and does not prevent these silent errors. However, these cases are rare (§ 7.2.1), and Kishu provides a blocklist file for users to force fallback recomputation for Co-variables containing objects belonging to these classes.
Alternative Delta Detection Methods
While Kishu’s VarGraphs can generalizably detect Co-Variable updates, there are specific cases in which more efficient detection methods are applicable, for example, (1) hashing (e.g., XXH64 (xxHash, 2023)) for array-like data (Team, 2024a)) and (2) rule-based static cell (e.g., df.head) identification for skipping update detection (§ 2.2). Kishu currently uses hash-based detection for some common array-likes (e.g., TensorFlow’s tensors (TensorFlow, 2024a)), but can be extended to incorporate (1) other efficient detection methods for specific data classes and (2) rule-based cell analyses.
7. Experimental Evaluation
In this section, we empirically study the effectiveness of Kishu’s time-traveling. We make the following claims:
  1. 1.
    Generalizable and Robust Mechanism: Kishu can identify modifications to, and correctly restore session states containing 146 object classes from common Data Science libraries. (§ 7.2)
  2. 2.
    Low Checkpoint Storage Cost: Kishu’s optimizations result in its per-cell-execution checkpoints being up to 4.55$\times$ smaller compared to those from the next best mechanism. (§ 7.3)
  3. 3.
    Low Checkpoint Times: Kishu’s incremental per-cell-execution checkpoints are created in up to 5.12$\times$ less time compared to the next best mechanism (§ 7.4)
  4. 4.
    Fast Incremental Checkout: Kishu’s novel incremental restoration is crucial to its sub-second checkout times — up to 8.18$\times$ and 4.18$\times$ faster than the next best mechanism for undoing cell executions and switching branches, respectively. (§ 7.5)
  5. 5.
    Low Overhead Delta Detection: Kishu incurs negligible runtime overheads on data science notebooks for capturing the state delta — less than 3.0% of the notebook session runtime and up to 4.08$\times$ less than alternative tracking approaches. (§ 7.6)
7.1. Experiment Setup
Datasets
We select 8 data science notebooks from Kaggle Grandmaster-level users or Github-hosted tool tutorials (e.g., Ray) (LABEL:tbl:workload), each featuring a popular data science library, which we categorize as in-progress (3/8) or final (5/8), with the former containing out-of-order cell executions and the latter lacking them. Notably, the final notebooks are also cleaned and contain memory/runtime optimizations. We empirically verify all notebooks follow traits discussed in § 2.2, and provide more details on our categorization (and its implications) in the appendix (§ 10).
We also select 146 data science library classes (summarized in LABEL:tbl:library) commonly used in data science notebooks, on which we evaluate Kishu’s correctness and robustness.
Methods
We evaluate Kishu against existing tools capable of enabling time-travelling on notebooks to various degrees:
  • CRIU (CRIU, 2023): Performs a system-level memory dump of the process hosting the notebook session. The session state is restored by loading the memory dump and reviving the process.
  • CRIU-Incremental (CRIU, 2023): CRIU with snapshot deduplication, storing only dirty memory pages in subsequent snapshots.
  • DumpSession (Foundation, 2023h): An application-level checkpointing tool that serializes the entire session state into one single file.
  • ElasticNotebook (Li et al., 2023a): An application-level notebook migration tool that balances data serialization and cell recomputation to achieve optimized session replication times.
  • Kishu+Det-replay: A checkpoint-optimized version of Kishu based on operation replay (Mohan et al., 1992; Netezza, 2024) that uses manual annotation [^6] Automatically detecting non-/determinism in executions is out of scope for this work. to skip incremental checkpointing for deterministic cells. These deterministic cells are replayed as necessary on checkout.
Ablation Study
We additionally compare the overhead of Kishu’s update detection mechanism with these tracking methods (§ 7.6):
  • IPyFlow (Shankar et al., 2022): A hybrid dynamic-static (i.e., AST analysis with live symbol resolution) for obtaining sub-variable (i.e., symbols, e.g., ls[x]) level granularity to perform reactive cell executions.
  • AblatedKishu (Check all): Always perform update detection for all Co-variables in the session state after each cell execution, regardless of whether they were accessed in the previous cell.
Methodology and Measurement
We run notebook cells sequentially and checkpoint after each cell execution. We checkout into the same state for Kishu and Kishu+Det-replay and into a fresh kernel/process for other methods. We measure the (1) checkpoint time (including both tracking and data writing) after each cell execution, (2) checkout time to restore the state from checkpoint files, and (3) tracking overhead spent by Kishu after each cell execution to track updates. We clear the page cache between runs.
Environment and Reproducibility
Experiments are performed on an Ubuntu server with 2 AMD EPYC 7552 48-Core Processors and 1TB RAM. All checkpoints are written to a mounted NFS, with disk read and write speeds of 519.8 MB/s and 358.9 MB/s, respectively. Our Github repository [^7] https://github.com/illinoisdata/kishu-vldb contains our implementation of Kishu, experiment notebooks/library classes, and scripts.
7.2. Generalized and Robust Time Traveling
This section compares the robustness of Kishu’s time-traveling to existing methods. We attempt to checkpoint and checkout session states containing objects from the 146 data science library classes and compare number of classes each method fails to checkout.
We report results in LABEL:fig:experiment_robust. Kishu completes time-traveling for all 146 libraries, handling 6 classes with multiprocessing and/or off-CPU data and 7 unserializable classes that CRIU and DumpSession fail on, respectively: unlike CRIU, Kishu utilizes reductions (§ 6.1) to store Co-variables, hence it can store distributed or off-CPU data (e.g., Ray’s dataset(Team, 2024b) or on-GPU tensors(TensorFlow, 2024b; PyTorch, 2024)) and unlike DumpSession, Kishu’s fallback recomputation allows it to restore Co-variables with (1) unserializable objects (e.g., pl.LazyFrame(Polars, 2024b)) or (2) serializable objects that can’t deserialize (e.g., bokeh.figure(Bokeh, 2024)). LABEL:tbl:experiment_robust_example_table summarizes these noteworthy classes.
7.2.1. Accurate Delta Detection
We verify Kishu’s delta detection accuracy by comparing two VarGraphs generated for each class object before and after (1) updating a class attribute (e.g., model.key = ’A’) or (2) updating nothing. We count the number of VarGraph differences for case (1) as successes and case (2) as false positives.
We report results in LABEL:tbl:experiment_fast_capture_accuracy. Kishu’s VarGraphs accurately captures object updates in 120 classes. While Kishu reports false positives in 14 classes, (e.g., due to dynamically generated reachable objects), they only affect Kishu’s efficiency (i.e., during time-traveling); however, Kishu maintains accuracy by considering these objects to be updated on access. We also find that 12 classes contain silent pickling errors (§ 6.2); nevertheless, Kishu reports these objects to be updated on access similar to false positives, and users may force their (fallback) recomputation if needed (§ 6.2). Notably, Kishu has no false negatives: Kishu will always report if an object is changed.
7.3. Small Incremental Checkpoint Sizes
This section compares Kishu’s checkpoint sizes with those of existing tools: we checkpoint the session state after each cell execution with each method and measure the total storage size of checkpoints.
We report results in LABEL:fig:experiment_cheap_checkpoint. Kishu’s cumulative checkpoint size is consistently the smallest (expectedly except for Kishu+Det-replay, explained shortly) and is up to 4.55$\times$ smaller than the next best alternative (HW-LM). ElasticNotebook, while the next best method on 6/8 notebooks and also has fault-tolerant mechanisms to checkpoint all 8 notebooks, can fall short in checkpointing time (§ 4.3). CRIU-Incremental, while also incrementally checkpointing, is not the next best method on any notebook, losing to ElasticNotebook and DumpSession on 6 and failing to checkpoint on 2 as it (1) incrementally checkpoints at the coarser memory page level (§ 2.4), and (2) does not handle off-CPU data and multiprocessing (§ 7.2). DumpSession fails on Qiskit as it cannot handle unserializable data, and CRIU incurs prohibitive storage costs (94GB, TPS) as it non-incrementally checkpoints at OS-level. While Kishu+Det-replay can save checkpoint storage cost of up to 3.95$\times$ vs. Kishu (StoreSales) by skipping checkpointing after deterministic cells, it (1) needs manual annotation and (2) can result in unacceptable checkout times (§ 7.5).
7.4. Low Incremental Checkpoint Time
This section compares the checkpoint time of Kishu with that of existing tools: we measure the total time spent by each method creating checkpoints after each cell execution.
We report results in LABEL:fig:experiment_fast_checkpoint. Kishu’s cumulative checkpointing time is the lowest (except for Kishu+Det-replay) on 5/8 notebooks, being only up to 15.5% of notebook runtime (HW-LM) and up to 5.12$\times$ faster (HW-LM) than the next best alternative on these notebooks. While CRIU-Incremental checkpoints faster than Kishu on 3/8 notebooks owing to memory dumping being faster than serialization for unit data, the difference is negligible (up to 3.03$\times$, StoreSales) compared to the reliability issues (§ 7.2), space inefficiency (§ 7.3), and slow checkout times (§ 7.5). Compared to ElasticNotebook, Kishu’s checkpointing is EAFP-based (Goodger, 2014): if it fails to store a Co-variable, it will simply recompute it upon checkout via fallback recomputation. This allows it to skip the profiling (i.e., for data sizes and serializability) required for ElasticNotebook’s optimization (for what to store/recompute), which causes checkpoint times slower than DumpSession on 2/8 notebooks.
7.5. Fast Incremental Checkout
This section compares the efficiency of Kishu’s incremental checkout with the (non-incremental) checkout of existing methods. We generate per-cell-execution checkpoints on the notebooks following the methodology in § 7.3 and § 7.4, then measure the time it takes for each method to checkout to a previous state (i.e. undo, § 7.5.1) or checkout to a different execution branch (§ 7.5.2).
7.5.1. Fast Execution Undo
For each notebook, we measure the time it takes to undo various dataframe and plot operation cells.
We report the results in LABEL:fig:experiment_checkout_undo. Kishu is the only method capable of incrementally checking out: it achieves sub-second cell execution rollbacks on all test cases, and is up to 8.18$\times$ faster than the next best alternative (StoreSales). While CRIU-Incremental achieves checkpoint times comparable with Kishu, it is up to 36$\times$ slower for checking out (StoreSales) and the slowest method for undos on 5/6 notebooks, due to it needing to piece together the memory snapshot of the notebook process to restore from multiple (incremental) checkpoint files. CRIU, DumpSession, and ElasticNotebook cannot incrementally checkout hence they cannot consistently perform sub-second undos. For example, the Sklearn notebook test case drops a column in an auxiliary dataframe that is 1.4MB in size (vs. the 133MB main dataframe). Kishu identifies that it only needs to load the auxiliary dataframe from before the cell execution and undoes the operation in 0.4 seconds; however, other non-Kishu+Det-replay methods all require the entire session state to be overwritten with a complete load of checkpoint data, taking an upwards of 6 seconds to do so (and in CRIU and CRIU-Incremental’s case, also killing and restarting the current notebook process).
7.5.2. Fast Path Exploration
For each notebook, we (1) run the notebook end-to-end, (2) checkout to the state before any models are trained, (3) rerun to the end of the notebook (thus creating a second branch), and measure the time taken to switch back to the first branch containing different models and plots.
We report the results in LABEL:fig:experiment_checkout_branch. Similar to § 7.5.1, Kishu performs sub-second branch switching on 4/6 notebooks by updating (only) models and plots differing between branches (i.e., not the input dataframes) and does so up to 4.18$\times$ faster than the next best alternative (StoreSales). While there is considerable divergence between branches in the StoreSales test case (i.e., new auxiliary dataframes are created along ML models and plots), Kishu still performs branch switching at a fast 1.73 seconds, which is 4.18$\times$ faster than the next method (DumpSession). While Kishu+Det-replay can potentially be faster than Kishu (TorchGPU) by replaying cells that allow it to bypass expensive data loading, it can also cause unacceptable checkout times (1050s, Cluster, from replaying an entire deterministic model fitting sequence); hence, cost-based optimization is required for Kishu+Det-replay to function, which we leave to future work.
7.6. Fast Delta Detection
This section investigates Kishu’s Co-variable granularity state tracking overhead by comparing the time taken by Kishu to track per-cell execution state delta with other tracking methods.
Cumulative Tracking Overhead (LABEL:tbl:experiment_fast_capture)
Kishu is consistently fastest at detecting state delta and is (1) up to 11.42$\times$ faster than the best out of IPyFlow and AblatedKishu (Check all) (HW-LM), and (2) only up to maximum of 2.03% of notebook runtime (Sklearn).
Per-Cell Tracking Overhead (LABEL:fig:experiment_fast_capture_percentage)
[…]
Notably, there exists further optimization opportunities for Kishu such as for (1) cells updating Co-variables with nested VarGraphs (e.g. list of strings text_neg, Sklearn cell 41) and (2) read-only printing cells (e.g., y_train[:10], HW-LM cell 67). Kishu incurs significant overhead on these cases (260$\times$ and 1.06$\times$ respectively; while they are of low absolute value due to the short cell runtimes (547ms/2ms and 2ms/2ms respectively), this indicates need for more efficient delta detection methods such as list hashing and rule-based detection, respectively (§ 6.2), which we leave for future exploration.
7.7. Workload Study
This section studies Kishu’s performance versus parameter sweeps on the degree of shared referencing between variables (i.e., Co-Variable size) (§ 7.7.1) and number of cell executions (§ 7.7.2).
7.7.1. Performance vs. Shared Referencing
We insert ten 64MB numpy arrays into a list (of state data in a Co-Variable). We evaluate Kishu’s checkpoint/checkout costs on cells modifying only one array in the list/Co-Variable. DumpSession’s and CRIU-Incremental’s performances are provided as comparison.
We report results in LABEL:fig:experiment_covariable. As Kishu detects updates and checkpoints at the Co-variable granularity, it checks for updates and then checkpoints all arrays in the list after each test cell. Kishu performs best when the list contains a low % of data in the state, as it can (1) limit the scope of update checking and (2) time-travel by saving/loading a small amount of state data. Kishu’s performance drops as the list Co-variable bundles the changed array with more unchanged data and is equivalent to DumpSession  [^9] Different serialization libraries cause different checkpoint/checkout time cost; there is also overhead with managing data with blob storage vs. writing directly to file (§ 6.1). when all 10 arrays (i.e., all data in the state) form one large Co-Variable: Kishu has to (1) check the whole state for updates and (2) save/update all data for checkpoint/checkout, while CRIU-Incremental can still checkpoint only the one changed array in the (640MB) Co-variable.
However, as our evaluation workloads suggest (LABEL:tbl:experiment_covariable_count), states typically consist of a large number of small Co-Variables (each containing 2.57% of state data on average, black vertical lines in LABEL:fig:experiment_covariable). For these typical cases, our Co-Variable-based approach significantly outperforms other baselines (§ 7.3, § 7.4).
7.7.2. Scalability to Long Notebook Sessions
We choose 2 Data Visualization notebooks (HW-LM and Qiskit, and randomly re-execute up to 1000 cells. [^10] This is the length of the longest observed notebook on Kaggle (Google and X, 2022), and $\sim$10 times the 97th percentile cell execution count of a notebook workload (Rule et al., 2018). We measure the (1) Checkpoint Graph size and (2) time to compute state difference for undoing 0-1000 cells (i.e., to a prior state) from the state after the 1000th cell execution.
We report results in LABEL:fig:experiment_scalability. The size of Kishu’s Checkpoint Graph scales linearly with the number of executed cells and is up to only 9MB at 1000 cells (LABEL:fig:experiment_scalability_a, HW-LM). Kishu’s time to compute state difference between current and target states scales linearly with the sum of cell count in the two states, up to 81ms for any checkout operation in a 1000 cell execution-long session. [^11] This is because Kishu traverses the Checkpoint Graph to find the lowest common ancestor state on checkout with an off-the-shelf algorithm (Wikipedia, 2024). These overheads are negligible on a more typical notebook—only 133KB and 5.8ms, respectively, on our longest experiment notebook (Qiskit, 85 cells).
8. Related Work
This section covers related work in PITR/checkpointing in various database applications (§ 8.1) and related notebook systems and their employed techniques for lineage tracing (§ 8.2). , and other applicable tools for saving/loading data in notebook states (§ 8.3)
8.1. PITR and Incremental Checkpointing
PITR in Relational Databases
In many DBMSs, mechanisms like ARIES (Mohan et al., 1992; Mohan and Rothermel, 1988) and its variants (Arulraj et al., 2016; Zheng et al., 2014; Soroush and Balazinska, 2013; Schule et al., 2019; Gray et al., 1981; Salem and Garcia-Molina, 1989; Verhofstad, 1978; Morrey and Grunwald, 2003) achieve durability and Point-in-time-Recovery (PITR) by combining (incremental) checkpointing and physical/operation logging (Oracle, 2024a; Group, 1996b; Soroush and Balazinska, 2013; Schule et al., 2019; Xu et al., 2017; Antonopoulos et al., 2019; Verbitski et al., 2017; Zheng et al., 2014). Persisting dirty objects (e.g., rows in relations) can enhance recovery efficiency by reducing the number of log entries to replay. ARIES (Mohan et al., 1992) identifies dirty objects at the page level by recording a RecLSN, i.e., the earliest modification time, in the dirty page table. This is possible because all the pages are controlled by the transaction system’s buffer manager. Kishu shares similarities: updates are periodically flushed. However, the core question is how to define those updates in computational notebooks. There are no buffer managers; moreover, variables/objects are dependent with inter-object references (i.e., memory pointers). Simply storing/restoring some variables independently thus can invalidate the state. Kishu addresses this with Co-Variables, atomic units of persistence that can correctly preserve inter-object dependencies. Kishu can still benefit from other PITR optimizations, e.g., non-blocking checkpointing/restoration (Mohan et al., 1992) and cost-based deterministic replay (Netezza, 2024) (§ 7.5.2), which we leave as future work.
PITR in Blob Storages
Blob storages have incorporated PITR for recovering its state to earlier points (Netezza, 2024; Group, 1996a; Azure, 2024a, b; Amazon, 2024b, a; LITEREPLICA, 2024). These systems typically log timestamped updates (e.g., Azure’s change feeds (Azure, 2024a) and SQLite’s binary log (LITEREPLICA, 2024)), which can be used to return the storage state to a previous snapshot. Computational notebooks require different approaches to determining state deltas, and Kishu addresses the unique challenge through Co-Variables and related techniques, which is the core novelty of this work.
Incremental and Differential Checkpointing in HPC
HPC and stream processing systems implement efficient incremental checkpointing (Lin et al., 2020; Keller and Bautista-Gomez, 2019; Bronevetsky et al., 2008; Apache, 2024). These works focus on high-frequency (e.g., ¿1M/second) updates to relatively simple data structures (e.g., KV-stores (Lin et al., 2020; Apache, 2024) and datasets (Keller and Bautista-Gomez, 2019; Bronevetsky et al., 2008)). For enhanced efficiency, log truncation (Lin et al., 2020) and copying (Keller and Bautista-Gomez, 2019) are employed to ensure the bounded size and robustness of incremental logs. Computational notebooks require different delta detection methods as notebooks involve more complex objects, e.g., graphs, tensors, dataframes (§ 2.2). While orthogonal, the existing optimization techniques for HPC may offer significant performance benefits if the core techniques for Kishu are expanded to handle a significantly larger number of executions (e.g., scripts) compared to interactive notebooks.
8.2. Notebook Systems and Techniques
Systems for Speeding up Data Exploration
There are a variety of works for enhancing data exploration efficiency in notebook-based systems (Wu et al., 2020; Koop and Patel, 2017; Shankar et al., 2022; Lee et al., 2021; Li et al., 2023b; Bäuerle et al., 2022; Guo and Seltzer, 2012; Brachmann et al., 2019). Reactive execution engines (Wu et al., 2020; Koop and Patel, 2017; Shankar et al., 2022) track cell reruns and rerun their dependent cells reactively to enforce consistent cell outputs. Notebook recommender systems (Lee et al., 2021; Li et al., 2023b) compute next-cell recommendations based on the current workflow. Symphony (Bäuerle et al., 2022) and B2 (Wu et al., 2020) enable point-and-click interactions with ML models and dataframes, respectively, by translation to equivalent code operations. Burrito (Guo and Seltzer, 2012) and Vizier (Brachmann et al., 2019) construct and visualize a history graph for versioning multi-language (Python, R, SQL) data science code. Diff-in-the-loop (Wang et al., 2022) enables graphical comparisons of dataframes. Kishu facilitates data exploration through efficient time-traveling, and thus, is orthogonal and complementary to these works.
Notebook State Versioning
Notebook state versioning has been explored by ForkIt (Weinman et al., 2021), which performs backtracking and branching to/from states by saving entire states with pickle (Foundation, 2023f). Our work versions notebook states in a much more efficient and robust manner compared to Forkit’s approach (which is equivalent to our DumpSession baseline in § 7) through efficient delta detection and incremental checkpointing and checkout.
Lineage Tracing in Notebooks
Lineage tracing aims to capture code dependencies, i.e., accessed and updated data of cells, and has been widely used in notebook systems for a variety of downstream tasks (e.g., reactive execution) (Manne et al., 2022; Macke et al., 2020; Ahmad et al., 2022; Li et al., 2023a; Cunha et al., 2021; Wu et al., 2020; Koop and Patel, 2017; Shankar et al., 2022; Lee et al., 2021; Li et al., 2023b; Brachmann et al., 2019; Deo et al., 2022). Tracing methods can be divided into (1) static code analysis using methods like AST decomposition (Foundation, 2023a), and (2) live code instrumentation resolving variable/data references at runtime. They are relatively (1) cheap but conservative, and (2) accurate but expensive, respectively. Lineage tracers often combine static and live analysis, to mitigate conservative assumptions of static analysis (Li et al., 2023a) with dynamically computed ’ground truths’ (Macke et al., 2020; Shankar et al., 2022; Deo et al., 2022). Unfortunately, to the best of our knowledge, all of these works, except for ElasticNotebook (Li et al., 2023a), detect modifications at the variable level, incorrectly disregard shared references between variables. Kishu’s innovations in lineage tracing is efficient live analysis through (1) modeling the state at a coarser Co-variable granularity and (2) quickly pruning update candidates. This allows Kishu to achieve low false positives (§ 7.2) with low overhead (§ 7.6).
8.3. Checkpointing Notebook Objects
Data Serialization for Checkpoint/Restore
Data in IPython-based (e.g., Jupyter) notebook session states can be saved with serialization libraries (Foundation, 2023f, g, e, d; MongoDB, 2023; MessagePack, 2024; DuVall, 2015; cloudpipe, 2024), on which a variety of existing checkpointing tools are built: On-disk KV-stores save individual variables (Team, 2023c; shove, 2024; 0xnurl, 2024; Python, 2024; Foundation, 2024b; chest, 2024), DumpSession (Foundation, 2023h) saves the session state in bulk, and ElasticNotebook (Li et al., 2023a) combines data storage/loading with cell replay for optimized session replication. These works do not checkpoint nor restore incrementally, or have limitations/require significant user effort: Dill and ElasticNotebook’s checkpoint files must be entirely loaded for restoration; while KV stores can store and load parts of a state, delta detection and shared reference preservation must be manually handled. In comparison, Kishu can perform low-overhead incremental checkpointing and checkout while preserving shared references (i.e., correctness) and works with almost all data science libraries (§ 7.2).
Memory Snapshotting
There exist OS-level checkpointing tools that can incrementally checkpoint a process for later restoration (CRIU, 2023; Ansel et al., 2009; Chen et al., 1997; Plank et al., 1994; Garg et al., 2018; Jain and Cooperman, 2020; Li and Lan, 2010). These tools identify and store dirty memory pages, then piece together the process image for restoration (CRIU, 2023): they often (1) cause large checkpoint sizes due to coarse page-level deltas (Elnozahy et al., 2002), (2) can only handle single processes (CRIU, 2024), and (3) can only restore from scratch: while we found a patent (Neary et al., 2007) and paper (Ferreira et al., 2011) enabling incremental checkpointing for multiprocessing jobs and with sub-memory-page granularity, respectively, we could not locate working implementations. In comparison, Kishu achieves lower checkpoint overheads via the logical Co-variable granularity deltas (§ 7.3), checkpoint multiprocessing notebooks via application-level instructions (§ 7.2), and incrementally restores via state difference detection and using existing kernel data (§ 7.5).
9. Conclusion
We have proposed Kishu, a new computational notebook system that offers efficient and fault-tolerant time-traveling between notebook states. Kishu captures session state evolution at a novel Co-variable granularity for efficient incremental checkpointing of state deltas, which Kishu then uses to perform incremental checkout with minimal data loading. Kishu’s contributions include (1) low-overhead state delta detection, (2) branch-based state versioning, and (3) generalizable time-traveling—preserving inter-variable dependencies and handling missing data with fallback recomputation. We have shown that Kishu is compatible with 146 data science object classes and reduces checkpoint storage size and checkout time by up to 4.55$\times$ and 8.18$\times$, respectively, on real-world notebooks.